OpenStudy (anonymous):

Please help!!!! Which is the following includes multiples of 12? a. 12,24,36.48 b. 6. 12,24,30 c. 12, 36, 60, 96 d. 24, 48, 54, 60

OpenStudy (anonymous):

a

OpenStudy (anonymous):

what is wrong with c. 12, 36, 60, 96 ?

OpenStudy (anonymous):

oh a and c, 12/12=1 , 24/12=2, 36/12=3, 48/12=4 60/12=5, 96/12 =8 so they are all multiples of 12

OpenStudy (anonymous):

yes...both of c and a....but,must be choose juz one answer???

OpenStudy (anonymous):

yes that is why I was confused!

OpenStudy (anonymous):

hmmm...its myb type error...where do you get this question???books???

OpenStudy (anonymous):

Weird question.. All of them include multiples of 12 heh.

OpenStudy (anonymous):

yes from book

OpenStudy (anonymous):

i agree with cazil....how can i help u...i think is type error...the answer is a and c